OpenStudy (anonymous):

\sqrt{18} - \sqrt{50}

OpenStudy (anonymous):

Tiana, the question is? :)

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{18} - \sqrt{50}\] \[\large \sqrt{9*2} - \sqrt{25*2}\] \[\large \sqrt{9}*\sqrt{2} - \sqrt{25}*\sqrt{2}\] \[\large 3*\sqrt{2} - 5*\sqrt{2}\] \[\large (3 - 5)*\sqrt{2}\] \[\large -2\sqrt{2}\] ======================================================= Therefore, \[\large \sqrt{18} - \sqrt{50} = -2\sqrt{2}\]

OpenStudy (jdoe0001):